OpenStudy (anonymous):

t^2+6t-7=(t+7)(t-1) Can you simplify?

OpenStudy (hba):

@boompolis You should start working on it and i will help you where you go wrong.

OpenStudy (anonymous):

t^2 + 6t / t-1 - 7 / t-1 t^2 + 6t - 7 / t-1 factoring t^2 + 6t - 7 = (t+7)(t-1) ... now simplify

OpenStudy (theviper):

\(\large{t^2+6t-7=(t+7)(t-1)}\) \(\large{\implies t^2+6t-7=t^2-t+7t-7}\) \(\large{\implies t^2+6t-7=t^2+6t-7}\) Therefore it is true means right :)\[\Huge{\color{red}{\checkmark}}\]

OpenStudy (anonymous):

so which one is the simplified answer

OpenStudy (anonymous):

@TheViper

OpenStudy (theviper):

\[\Huge{\color{red}{t^2+6t-7}}\]

OpenStudy (anonymous):

thanx

OpenStudy (theviper):

ur welcome :)